A rectangular field with area 4500 square meters is enclosed with 300 meters of fencing. Find its dimensions. (HINT: set up a system)
Answer by mananth(16949)

let length be l
width w
..
Area = l*w
l*w= 4500
l= 4500/w
...
perimeter = 2(l+w)
2(l+w)=300
/2
l+w = 150
..
plug value of l
4500/w +w = 150
(4500+w^2)/w=150
4500+w^2=150w
w^2-150+4500=0
discriminant = 4500
solve for x using quadratic equation
x1=(150+sqrt(4500))/2
x1=108.5 m
x2= 150-sqrt(4500))/2
x2= 41.5
The dimensions are 108.5m by 41.5 m

...
m.ananth@hotmail.ca